Nerdyhuman:

Hey, I have a question with 2 parts that I am not sure how to do. Thank you!

Sailor:

Is there a screenshot of it or something?

Rylee88:

What's the Question(s), So we can help you.

ILOVESPAGHETTI:

Is there a image or something you can provide?

Rylee88:

There offline.

Sailor:

@rylee88 wrote:
There offline.
Wow. Just... WOW

ILOVESPAGHETTI:

@sailor wrote:
@rylee88 wrote:
There offline.
Wow. Just... WOW
dang how are we supposed to answer LMAO

Sailor:

@ilovespaghetti wrote:
@sailor wrote:
@rylee88 wrote:
There offline.
Wow. Just... WOW
dang how are we supposed to answer LMAO
ikr

Nerdyhuman:

A 3.0 cm lighted candle is placed 20 cm from a concave spherical mirror with a radius of curvature of 30 cm. (a) Where should a screen be placed (in cm) in order to see the candle's image clearly? (Give your answer as a distance from the center of the mirror.)

Nerdyhuman:

sorry I had to handle something. I am back online

Sailor:

Ok

ILOVESPAGHETTI:

@nerdyhuman wrote:
sorry I had to handle something. I am back online
It's okay

Nerdyhuman:

so what do I do for the problem

Nerdyhuman:

@vocaloid

Florisalreadytaken:

A 3.0 cm lighted candle is placed 20 cm from a concave spherical mirror with a radius of curvature of 30 cm. (a) Where should a screen be placed (in cm) in order to see the candle's image clearly? (Give your answer as a distance from the centre of the mirror.) So, were given that the mirror is a concave spherical mirror, which means that the formula for it would be \( \frac{1}{F}=\frac{1}{U_{\ell}}+\frac{1}{V} \) where: \(\large\mathcal{r} _{i}\) \(= -30cm\) \( U_{\ell} =-20cm\) \( F=\frac{r_{c}}{2}=-\frac{30cm}{2}=-15cm \) plugging what we have into the formula we would get: \( - \frac{1}{15cm}=-\frac{1}{ 20cm }+\frac{1}{V} \) find the LCM: \( -\frac{1}{15cm}\times\:60v=-\frac{1}{20cm}\times\:60v+\frac{1}{v}\cdot \:60v \) \( -4v(cm)=-3v(cm)+60 \) \( -4v(cm)+3v(cm)=60 \) \( -v(cm)=60 \) \( \boxed{ v=-60(cm) }\) Looking at the answer, we can say that the screen should be settled \( 60cm \) AWAY from the centre of the candle to see a clear image of it.

Nerdyhuman:

thank you What is the minimum height (in cm) of the screen in order to see the candle's complete image? what about this?

Florisalreadytaken:

minimum height (in cm) of the screen in order to see the candle's complete image basically means the magnification. the formula for that is: \( \mathrm{m_{ag}}=\frac{\mathrm{h}^{\prime}}{\mathrm{h}}=-\frac{\mathrm{v}}{\mathrm{u}} \) can you work out the answer?

Nerdyhuman:

so is 60 the v?

Nerdyhuman:

trying to figure out all of the letters lol but then it will be easy to plug in

Florisalreadytaken:

its always NEGATIVE. so \( V=\color{red}{-}60cm \)

Nerdyhuman:

what is u?

Nerdyhuman:

like what does it represent

Florisalreadytaken:

That's a bit rude for me to solve a problem, and the asker not to know what I even did. \(U\) is the distance ( \(\ell\) ) between the object and the mirror.

Nerdyhuman:

oh nevermind I see it now sorry I think I may have it

Nerdyhuman:

okay so it would be -60/20 correct? Isn't h'/h supposed to equal the -60/20 as well? If that's the case, I have 3 as an answer but the program says that answer is wrong

Florisalreadytaken:

I reckon it's supposed to be negative 3cm.

Nerdyhuman:

I tried that as well

Nerdyhuman:

maybe try 2 and put the 30 in the 20s place?

Nerdyhuman:

It isn't 2 either

Nerdyhuman:

I figured it out. It's 9 because we are figuring out the screen height. Both sides must equals 3.

Florisalreadytaken:

sweetas!